# Learning Naturally Aggregated Appearance for Efficient 3D Editing

### v1

@@ 3.3 Canonical Projection Initialization
where $\mathbf{x}\in\mathbb{R}^{3}$ is a 3D point in Euclidean space. This design transforms the points in such a way that they are distributed proportionally to the disparity in a unit sphere. Accordingly, the voxel grid $\phi_{G}$ is a cube with range $[-1,1]^{3}$.
Using the contracted representation in Eq. 8, we have $(p_{x^{\prime}},p_{y^{\prime}},p_{z^{\prime}})\triangleq f_{c}(\mathbf{p}_{xyz})$. The canonical projection initialization $P_{c}$ from $\mathbf{p}_{xyz}$ in 3D space to the 2D canonical image pixel $(\tilde{p}_{u},\tilde{p}_{v})\triangleq\tilde{\mathbf{p}}_{uv}$ can be formulated as follows:
@@
3.4 Design and Regularization
Positional and hash encoding. The projection offset in  Eq. 4 employs Fourier positional encoding or multi-resolution hash encoding to capture high-frequency information. For $\gamma_{d}$, we specifically employ positional encoding [48]; as for $\gamma_{p}$, we choose either positional encoding or hash encoding [32]. The positional encoding is defined as $\gamma_{pe}(\cdot):\mathbb{R}^{3}\to\mathbb{R}^{3\times(1+2K)}$ to encode 3-dimensional vector $\mathbf{x}$ up to $K$ frequencies as $\gamma_{pe}(\mathbf{x})=[\mathbf{x},F_{1}(\mathbf{x}),...,F_{K}(\mathbf{x})]$, where $F_{k}(\mathbf{x})=[\sin(2^{k}\mathbf{x}),\cos(2^{k}\mathbf{x})]$; for hash encoding, we have $\gamma_{h}(\cdot):\mathbb{R}^{3}\to\mathbb{R}^{3+DK}$ to encode the vector $\mathbf{x}$ by a $K$-resolution hash grid with $D$-dimensional feature per layer as $\gamma_{h}(\mathbf{x})=[\mathbf{x},H_{1}(\mathbf{x}),...,H_{K}(\mathbf{x})]$, where $H_{k}(\mathbf{x})$ is a $D$-dimensional feature vector interpolated by $\mathbf{x}$ at the $k$-th resolution.
Anneal encoding. Motivated by Nerfies [38], the positional or hash encoding can incorporate an optional annealing learning strategy. To achieve this, we introduce a weight factor $w^{n}_{k}=\frac{1}{2}(1-\cos(\alpha^{n}_{k}\pi))$ for some encoded frequency $F_{k}^{n}$ or $H_{k}^{n}$ at some training step $n$, such that we have $F_{k}^{n}(\cdot)=w^{n}_{k}F_{k}(\cdot)$ and
$\alpha^{n}_{k}=\min(\max(\frac{n-N_{s}}{N_{e}-N_{s}}K-k,0.0),1.0),$ (10)
Progressive training. Similar to [3, 31, 47], we apply progressive scaling for our voxel grid $\phi_{G}$ and canonical image $\phi_{I}$ for a coarse-to-fine learning process. At specific scaling-up milestone steps, we increase the $\phi_{G}$ voxel count by a factor of 2 and the $\phi_{I}$ pixel count by a factor of 4.
$\mathcal{L}_{uv}=\left\|\Delta\mathbf{p}_{uv}\right\|_{2}^{2}.$ (11)
Optimization objective. The final optimization process of our method to model the scene for efficient editing can be formulated as follows:
Haque et al. [14]
Ours
Reference Image
Haque et al. [14]
Ours
Reference Image
Figure 4: Visual comparison of scene stylization results on the Instruct-NeRF2NeRF dataset given different text prompts.
@@ 3.4 Design and Regularization
Figure 6: Visual comparison of foreground and background extraction results on the LLFF dataset.
@@
Figure 7: More visualization of content extraction results in different complex settings on the LLFF dataset.
Edited Novel View 1
Edited Novel View 2
Edited Novel View 3
@@
Figure 8: Visualization of texture editing (i.e., drawing) results rendered at different novel viewpoints on the LLFF dataset.
4 Experiments
4.1 Experimental Setup
Datasets. The LLFF [30] dataset comprises 8 real-world scenes, where each scene is accompanied by several training images captured by handheld cameras placed in a rough grid pattern. We evaluate this dataset at a resolution of $1008\times 756$ after downscaling the images by a factor of 4.
The Replica [46] dataset is a collection of various high-quality and high-resolution 3D reconstructions of indoor scenes with clean and dense geometry. We evaluate the 14 panorama scenes as processed in SOMSI [13], where each scene is rendered as a grid of equally spaced spherical images at a resolution of $1024\times 512$.
The Instruct-NeRF2NeRF [14] dataset consists of 6 scenes captured in either an object-centric or forward-facing manner. We leverage this dataset specifically to model and edit forward-facing scenes that prominently feature human subjects for qualitative evaluation.
Implementation details. Our 3D editing pipeline involves a two-step process: training a per-scene reconstruction model and subsequent explicit edits on the canonical image $\phi_{I}$ for 3D scene editing. By default, we optimize a per-scene model for $60$k steps using the Adam optimizer [22] with an initial learning rate of $0.1$ for both the explicit 3D density grid $\phi_{G}$ and 2D canonical image $\phi_{I}$, and a learning rate of $0.001$ for the implicit projection field $P$ with learnable parameter $\phi_{P}$. All experiments are conducted and tested on a single RTX A6000 GPU. For further details and hyperparameters used in training different models, please refer to the supplementary materials.
Table 1: The reconstruction results on the LLFF dataset in terms of PSNR. Our PE-based models, which exhibit superior editing capacity, demonstrate satisfactory performance, while our hash-based models achieve reconstruction comparable to NeRF [31] and DVGO [47].
@@ 4.1 Experimental Setup
Table 2: The reconstruction results on the Replica dataset.
[…]
4.2 Evaluation on Editability
By performing explicit edits on the canonical image $\phi_{I}$, our model is capable of propagating the editing effects through the learned projection field $P$. This enables various 3D scene editing functionalities, such as scene stylization, context extraction, and texture editing. Specifically, except for explicit texture editing, we leverage several state-of-the-art models for 2D image editing. Specifically, we utilize the prompt-guided ControlNet [61] for scene stylization and Segment-anything (SAM) [23] for content extraction.
Scene stylization. We conduct a comparative analysis between our method and three state-of-the-art stylization methods: ARF [60], Ref-NPR [62], and Instruct-NeRF2NeRF [14]. All three baseline methods require additional optimization processes to achieve stylization with a specific style, somehow limiting their model practicability. Specifically, ARF and Ref-NPR rely on one or multiple stylized reference images, while Instruct-NeRF2NeRF utilizes text prompts through Diffusion models [4] for guidance.
In Fig. 3, we demonstrate some comparing visualizations evaluated on the LLFF dataset, where we can see that both ARF and Ref-NPR cannot edit the 3D scene to be as visually consistent as our method with respect to the given reference style image. This behavior can be attributed to ARF and Ref-NPR’s approach of adjusting NeRF model weights through loss functions to fit the given style, resulting in implicit control over the final edited style of the scene. In contrast, by editing the explicit canonical RGB image $\phi_{I}$ of our method, we can propagate the appearance to the 3D color field through the projection function $P$, enabling explicit appearance control for editing. We further showcase some visual results of the Instruct-NeRF2NeRF dataset using different text prompts for the third baseline in Fig. 4. Although both of our methods can successfully edit the scene into the desired style, optimizing the NeRF model using the Instruct-NeRF2NeRF method takes more than 10 hours to achieve stationary performance, while our method requires no extra re-optimization. Lastly, we present more visual editing results on the panorama Replica data in Fig. 5.
Content extraction. We evaluate our method with the state-of-the-art DFFs [24] method, which enables NeRFs to decompose a specific object given a text or image-patch query. In Fig. 6, we show some comparing visualization of foreground and background extraction. The evaluation of baseline DFFs is only based on text query according to its official codebase. As shown in Fig. 7, we demonstrate more visualization with various extraction goals, including multi-object extraction, custom-design content extraction, and complex object extraction, and we can even do stylization specifically on the extracted object.
Texture editing. We further present an additional application to do textural appearance editing of the scene by drawing or painting the canonical image. In Fig. 8, we can observe that our method ensures both appearance and 3D consistency in novel views.
4.3 Trade-off between Fidelity and Editability.
As a neural editing method without the need for re-optimization, we compensate reconstruction quality for editing capacity. In this section, we examine the reconstruction capacity of our method by trying different designs for learning the projection offset $P_{o}$, whose inputs are the query 3D position $\mathbf{p}_{xyz}$ and its viewing direction $\mathbf{d}$. Specifically, our findings indicate that utilizing a positional encoding (PE) for position leads to superior editing capacity, whereas models with hash encoding exhibit higher reconstruction performance compared to the PE-based models. The corresponding PSNR results are shown in Tabs. 1 and 2.
4.4 Ablation Study
Additionally, we conduct ablation studies to evaluate the effectiveness of different model components in terms of both editability and reconstruction fidelity. Specifically, we choose the trex scene in the LLFF dataset as our evaluation. The corresponding reconstruction statistics in terms of PSNR are shown in Tab. 3.
Explicitness and implicitness. Setting I and II replace the explicit representation of the density grid $\phi_{G}$ and the canonical image $\phi_{I}$ with implicit representations modeled by MLP networks, respectively, where their statistical result drop significantly. Moreover, in setting $1$, it fails to learn accurate geometries for reconstruction, and in setting $2$, it fails to learn a satisfying canonical image for editing purposes. Based on these findings, we decide to utilize explicit representations for both the density grid $\phi_{G}$ and the canonical image $\phi_{I}$.
Canonical projection initialization. Setting III ablates the canonical projection initialization $P_{c}$. The performance without $P_{c}$ is quite tricky, where the PE-based model experiences a drop while the hash-based model shows an increase. Nonetheless, both models lose the ability to perform editing tasks as the learned canonical images $\phi_{I}$ deviate from natural images to latent color maps.
Learnable projection offset. In Setting IV, the removal of view-dependence from the learnable projection offset $P_{o}$ leads to a minor decrease in performance, as the model no longer considers viewing directions. In Setting V, where the entire projection offset $P_{o}$ is eliminated, a significant drop in performance is observed. More visual analyses are shown in the supplementary materials.
Table 3: Ablation studies of model components on trex scene.
[…]
5 Discussion and Conclusion
The key to aggregating 3D appearance as an image for editing is to find a natural mapping from 3D points to 2D pixels. While such a mapping is non-trivial when it comes to 360-degree inward-facing scenes, and hence we mainly explore the potential of our method with forward-facing and panorama outward-facing data. As the field of neural scene representation progresses, we anticipate future work can successfully address this issue.
In summary, we propose AGAP, an editing-friendly and efficient solution for neural 3D scene editing. We leverage an explicit 2D appearance aggregation to replace the color field in NeRFs. Such a design allows users to directly perform 2D edits on images, eliminating the need for dealing with geometries or implicit fields, thus simplifying the editing process. Our approach showcases superior editing results on multiple datasets without any laborious re-optimization procedure.
Appendix
Column A
Column B
Column C
Column D
Column E
Column F
PE (I)
Hash (I)
PE (II)
Hash (II)
PE (III)
Hash (III)
PE (IV)
Hash (IV)
PE (V)
Hash (V)
PE (Full)
Hash (Full)
Canonical Image $\phi_{I}$
Canonical View
Canonical Depth
Canonical Image $\phi_{I}$
Canonical View
Canonical Depth
Figure A1: Visualization analysis of the ablation studies for both PE and hash models conducted on the trex scene in the LLFF dataset. Columns A and D are the learned canonical image $\phi_{I}$. Columns B and E are the rendered novel view images at the canonical position (global origin) with the canonical pose (identity rotation). Columns C and F are the corresponding depth maps of the rendered novel views. The zoomed-in patches for subfigure IV-A and Full-A are shown in Fig. A2.
IV: w.o. projection offset $P_{o}$
Full: w. projection offset $P_{o}$
Figure A2: Zoomed-in patches of the canonical images $\phi_{I}$ of the PE model with setting IV (subfigure IV-A) and our full PE model (subfigure Full-A). Without the projection offset, the canonical image fails to include the occluded background ceiling information. Our full model can handle occlusion and naturally aggregate the background ceiling inside the canonical image.
Appendix A Ablation Analysis
The ablation studies (settings I to V) presented in the main paper are conducted for both PE and hash models on the trex scene in the LLFF dataset. In this section, we supply the ablation studies with more corresponding visual analysis, offering more detailed insights. Specifically, we present the reconstruction PSNR statistics in Tab. A1, which match those reported in the main paper. For visual results, Fig. A1 shows the learned canonical image $\phi_{I}$ and the novel viewpoint image with the pseudo canonical camera pose plus its corresponding depth map for each ablation setting. The last row results in Tabs. A1 and A1 indicate that the PE model achieves a satisfactory reconstruction and a superior canonical image $\phi_{I}$ for explicit editing, whereas the hash model excels in terms of reconstruction quality.
Setting I: Replace the explicit density grid $\phi_{G}$ with an implicit one modeled by MLP networks. The learning rate for the replaced implicit density field is set to $10^{-3}$. We can see from Tab. A1 that the reconstruction qualities drop significantly as they fail to learn correct depths as shown by subfigure I-C and I-F in Fig. A1.
Setting II: Replace the explicit canonical image $\phi_{I}$ with an implicit one modeled by MLP networks. The learning rate for the replaced implicit canonical field is also configured to be $10^{-3}$. Based on the findings in Tab. A1, the reconstruction performance of both the PE and hash models experiences degradation of more than $3$ dB and $5$ dB, respectively. Further examination of row II in Fig. A1 reveals that while both models demonstrate satisfactory depth map learning capabilities, they exhibit limited capacities to generate high-quality canonical images for editing purposes.
Table A1: Ablation studies of model components on trex scene. Same as in the main paper.
[…]
Setting III: Remove the canonical project initialization $P_{c}$ in the projection field $P$. Although in Tab. A1 for setting III, we find that the PE model achieves satisfactory performance and the hash model even performs better than our full hash model in terms of PSNR, the learned canonical images (i.e., subfigure III-A and III-D in Fig. A1) are optimized to be latent color representations instead of natural images, resulting in a loss of editability.
Setting IV: Remove the project offset $P_{o}$ in the projection field $P$. Recall that the learned canonical image $\phi_{I}$ is determined by placing a pseudo canonical camera at the global origin. By comparing subfigure IV-A and Full-A of the PE models, we can see that without having the projection offset, the canonical image (subfigure IV-A) cannot “look through” the foreground ceiling, sharing similar behaviors as the canonical view (column B). However, by adding the projection offset, we enable the capacity of our method to handle occlusions. As demonstrated in subfigure Full-A, the final canonical image $\phi_{I}$ generated by our full models learns to naturally aggregate the occluded background ceiling within the image by leveraging the projection offsets from 3D points to 2D pixels. Please also refer to the corresponding zoomed-in patches in Fig. A2.
Setting V: Remove view dependency in the projection offset $P_{o}$ of the projection field $P$. The omission of viewing directions as the input for projection offset $P_{o}$, as observed from the last two rows in Tabs. A1 and A1, leads to a marginal decrease in reconstruction quality and yields slightly inferior canonical images.
Appendix B Training Details
Our 3D editing pipeline involves a two-step process: (1) we first train a per-scene reconstruction model using the proposed AGAP representation, which includes an explicit density grid $\phi_{G}$, an explicit canonical image $\phi_{I}$, and an associated projection field $\phi_{P}$; (2) we can then perform explicit 2D edits on the canonical image $\phi_{I}$ for 3D scene editing, including scene stylization, content extraction, and texture editing. All experiments, including training on various scenes from different datasets, are conducted and tested on a single RTX A6000 GPU, with specific hyperparameter details outlined in Tab. A2.
Table A2: Hyperparameters for training various scenes in the LLFF [30], Replica [46, 13], and InstructNeRF2NeRF (IN2N) [14] datasets.
[…]
@@ Appendix B Training Details
$\phi_{G}^{*},\phi_{I}^{*},\phi_{P}^{*}=\operatorname*{arg\,min}_{\phi_{G},\phi_{I},\phi_{P}}{\mathcal{L}_{color}+\mathcal{L}_{uv}+\mathcal{L}_{tv}},$ (A1)
where the second and the third terms are controlled by their corresponding weight factors $\lambda_{uv}$ and $\lambda_{tv}$, respectively. To be specific, the weight factor $\lambda_{uv}$ is set as $10^{-5}$ for forward-facing scene and $10^{-1}$ for panorama scenes; the weight factor $\lambda_{tv}$ is set as $10^{-5}$ for forward-facing scene and $10^{-4}$ for panorama scenes. Note that for panorama data, we set both the projection and total variation regularization to be relatively large because the depths of indoor scenes in the Replica dataset are sometimes ambiguous due to large walls in the background, where the total variation term is disabled after $20000$ steps to learn depths in details.
@@
At specific scaling-up milestone steps, we increase the $\phi_{G}$ voxel count by a factor of 2 and the $\phi_{I}$ pixel count by a factor of 4. For the forward-facing datasets (i.e., LLFF and InstructNeRF2NeRF), the voxel grid $\phi_{G}$ is scaled up at $\{2000,4000,6000,8000\}$ training steps and the canonical image $\phi_{I}$ is scaled up at $\{8000,16000\}$ training steps. For the panorama Replica dataset, the voxel grid $\phi_{G}$ is scaled up at $\{2000,4000,6000,8000,10000,12000,14000,16000\}$ training steps, and the canonical image $\phi_{I}$ is scaled up at $\{4000,8000,12000,16000\}$ training steps.
Size of voxel grid $\phi_{G}$ and canonical image $\phi_{I}$. After the progressive scaling up, The final resolution of the voxel grid $\phi_{G}$ is set as $384\times 384\times 256$ for forward-facing scenes and $320\times 320\times 320$ for panorama scenes.
In all the experiments, we set the final height $H_{I}$ of the learnable explicit canonical image $\phi_{I}$ as $768$. For panorama data, the canonical image width $W_{I}$ is set to be $1536$ according to the definition of Equirectangular projection. For forward-facing data, the canonical image width $W_{I}$ is adaptively calculated according to the width-height aspect ratio of the training images and the computed bounding box of the scene in NDC space. Denoting the bounding box in NDC space as $(x^{\prime}_{min},x^{\prime}_{max})$ in $x^{\prime}$ dimension, $(y^{\prime}_{min},y^{\prime}_{max})$ in $y^{\prime}$ dimension, and $(z^{\prime}_{min},z^{\prime}_{max})=(-1,1)$ in $z^{\prime}$ dimension and the aspect ratio as $r_{I}$, we can then calculate the canonical image width as:
$W_{I}=H_{I}\times r_{I}\times\frac{x^{\prime}_{max}-x^{\prime}_{min}}{y^{\prime}_{max}-y^{\prime}_{min}}.$ (A2)

### v2

@@ 4.1 Data Types
where $\mathbf{x}\in\mathbb{R}^{3}$ is a 3D point in Euclidean space. This design transforms the points in such a way that they are distributed proportionally to the disparity in a unit sphere. Accordingly, the voxel grid $\phi_{G}$ is defined as a cube with range $[-1,1]^{3}$. Given $f_{c}(\mathbf{p}_{xyz})\triangleq(p_{x^{\prime}},p_{y^{\prime}},p_{z^{\prime}})$, the canonical projection $P_{c}$ from a 3D point $\mathbf{p}_{xyz}$ to a 2D canonical image pixel $\tilde{\mathbf{p}}_{uv}\triangleq(\tilde{p}_{u},\tilde{p}_{v})$ can be formulated as follows:
@@
Figure 6: Visual comparison of foreground and background segmentation results on the LLFF dataset.
Object-centric data. The NeRF-Synthetic [46] dataset comprises synthetic objects with intricate geometry and realistic materials. Each object in the dataset includes 100 training views and 200 test views, all rendered at a resolution of $800\times 800$. The DTU [24] dataset is formulated in a object-level forward-facing setting, where training images are on quarter hemispheres ($\sim\frac{1}{8}$ spheres) with object masks. We evaluate this dataset at a resolution of $600\times 800$ after downscaling the images by a factor of 2. The NeUVF [41] dataset capture human head videos using 12 calibrated cameras located at a hemisphere with approximately 120. We use their first frames for static scenes, to specifically to model and edit the scenes that prominently feature human subjects for qualitative evaluation.
Analogous to Earth map, we can place the pseudo canonical camera at the global origin to model object-centric data using a (partial hemispheric) Equirectangular projection. Similar to Eq. 10, the canonical projection $P_{c}$ from a 3D point $\mathbf{p}_{xyz}\triangleq(p_{x},p_{y},p_{z})$ to a 2D canonical image pixel $\tilde{\mathbf{p}}_{uv}\triangleq(\tilde{p}_{u},\tilde{p}_{v})=(\tan^{-1}(\tfrac{p_{y}}{p_{x}}),\sin^{-1}(p_{z}))$.
Unbounded 360-degree data. The Mip-NeRF 360 [3] and the IN2N [18] dataset comprise unbounded outdoor and indoor scenes, which is the most challenging case. Each scene showcases a complex background along with a central object or area, captured at varying high resolutions.
Learning a good 3D-to-2D projection field $P$ for unbounded 360-degree scenes is indeed a non-trivial research challenge [12]. We leverage two canonical images for modeling the foreground central objects using (hemispheric) Equirectangular projection and the unbounded background by the contracted formulation as panorama, respectively.
Edited NVS 1
Edited NVS 2
Reference
Edited NVS 1
Edited NVS 2
Reference
Figure 7: Visualization of texture editing (i.e., drawing) results at different novel viewpoints on the LLFF, NeUVF, and DTU dataset.
Figure 8: A user study of our method with different alternatives in terms of both perceptual editing quality and consistency.
4.2 Implementation
Our pipeline involves a two-step process: 1) training a per-scene reconstruction model; 2) subsequent explicit edits on the canonical image $\phi_{I}$ for efficient 3D scene editing.
Training details. By default, we optimize a per-scene model for $60$k steps using the Adam optimizer [33] with an initial learning rate of $0.1$ for both the explicit density grid $\phi_{G}$ and the canonical image $\phi_{I}$, and a learning rate of $0.001$ for the implicit projection field $P$ with learnable parameters $\phi_{P}$. All experiments are conducted and tested on a single RTX A6000 GPU. For other implementation details and hyperparameters, please see the supplementary materials.
Editing pipeline. After we obtain the pre-trained model using our novel 3D representation, users can perform explicit edits as shown in Fig. 4 on the canonical image $\phi_{I}$ for various efficient 3D scene editing functionalities, such as scene stylization, instance segmentation, and texture editing. Our model can propagating the editing effects through the learned projection field $P$. Specifically, we utilize the prompt-guided ControlNet [85] for scene stylization and Segment-anything (SAM) [34] for instance segmentation. As for texture editing, users are free to draw or write on the canonical image.
Figure 9: More visualization of editing results on the 360 and object-centric datasets.
Table 2: Reconstructed PSNR on the LLFF and Replica datasets.
@@ 4.2 Implementation
LLFF [45] 24.13 -
NeRF [46] 26.50 -
DVGO [66] 26.34 -
SOMSI [17] - 39.54
Ours (PE) 24.83 38.42
Ours (Hash) 26.20 38.68
4.3 Evaluation on Editability
Scene stylization. We conduct a comparative analysis between our method and three state-of-the-art stylization methods: ARF [84], Ref-NPR [87], and IN2N [18]. Specifically, ARF and Ref-NPR rely on one or multiple stylized reference images, while IN2N utilizes text prompts through Diffusion models [4] for guidance. All these baseline methods require additional optimization processes to achieve stylization given a specific style, while our method is optimization-free. As shown in Tab. 1, our editing speed for stylization is approximately 20$\times$ faster than ARF and Ref-NPR and approximately 500$\times$ faster than IN2N.
Table 3: PSNR ablations of model components on trex scene.
@@ 4.3 Evaluation on Editability
I. No canonical projection $P_{c}$ 23.18 27.56
II. No projection offset $P_{o}$ 23.81 23.88
III. No viewdir $\mathbf{d}$ in projection offset $P_{o}$ 25.50 26.76
Full model 25.85 27.24
In Fig. 3, we demonstrate some comparing visualizations evaluated on the IN2N dataset and the LLFF dataset. In the first row, we present visual results from the IN2N dataset using different text prompts. While both our methods can effectively edit the scene to the desired style, the IN2N baseline necessitates approximately 3 hours to optimize the underlying NeRF model per edit, whereas our method requires no additional re-optimization. In the second row, we evaluate our method alongside the ARF and Ref-NPR baselines using the LLFF dataset. Both ARF and Ref-NPR produce implicit global stylization that looks like the reference style. However, our method achieves superior color and textural consistencies aligning with the provided reference style image. We present more visual results of stylization in Fig. 5 on the panorama dataset.
Instance segmentation. We evaluate our method with the state-of-the-art DFFs [35] method, which enables NeRFs to decompose a specific object given a text or image-patch query. The evaluation of baseline DFFs is only based on text query according to its official codebase. In Fig. 6, we show some comparing visualization of foreground and background segmentation. Our method allows users to do 3D instance segmentation easily by applying 2D segmentation of the desired objects in the canonical image.
Texture editing. Our method can further do textural appearance editing of the 3D scene by drawing or painting on the canonical image. In Fig. 7, we can observe that our method ensures both appearance and 3D consistency in novel views. For the first sample in row one, we paint an “AGAP” logo on the marble pedestal of the fern plant; for the second sample in row two, we draw two “ladybirds” on the leaves at the right bottom. We further present more visualizations in Fig. 9 and the supplementary materials.
User study. A user study comparing perceptual quality with baseline methods is presented in Fig. 8. We expect our approach to achieve comparable visual performance with existing alternatives, yet be far more efficient (at least 20$\times$ faster) and easier (e.g., supporting more types of editing) to use in practice as in Tab. 1. The user study involves 43 participants, each answering 10 comparison questions against baseline methods to select the best option:
  • For reference-based stylization (i.e., ARF and Ref-NPR), participants select the best aligned with the reference image in terms of editing quality and consistency.
  • For instruction-based stylization (i.e., IN2N and GaussianEditor), participants evaluate overall aesthetic appeal based on the provided instruction text prompt.
  • For segmentation comparisons (i.e., DFFs), participants determine whether ours or DFFs provides superior segmentation accuracy and editing performance.
(a) Ours (w/ $P_{c}$)
(b) Ours (w/o $P_{c}$)
(c) UV map [77]
Figure 10: Canonical image of ablating different 2D projections.
Stylization NVS
@@
Segmentation NVS
Figure 11: Hash models showcase a moderate level of editability.
4.4 Ablation and Analysis.
Editability. When designing the projection field, we find that having a good canonical projection is essential for natural and user-friendly editing. As presented by Setting I in Tab. 3, where the canonical projection is removed it while keeping the others unchanged, we can see that the PE model experiences a drop while the hash model shows an increase. However, both models lose the ability to perform editing tasks as the learned canonical images $\phi_{I}$ deviate from natural images to latent color maps, where the visual comparison is shown in Figs. 10(a) and 10(b). Apart from the various options for suitable canonical projection when handling different data types in Sec. 4.1, indeed, there exists some other projection functions like UV maps, but apparently they are hard to edit, as demonstrated in Fig. 10(c).
Reconstruction fidelity. AGAP, as a new efficient editing pipeline for 3D modeling, we also report the PSNR results here for completeness to showcase the capacity of faithful reconstruction in Tab. 2. We examine the capacity of our models with PE and Hash designs for the input $\mathbf{p}_{xyz}$ to learn the projection offset $P_{o}$. We find that PE models lead to superior editing capacity, whereas hash models prioritize reconstruction quality, meanwhile, maintaining a moderate level of editability as shown in Fig. 11.
Ours (w/o offset $P_{o}$)
Ours (w/ offset $P_{o}$)
Figure 12: Reconstruction and editing visual patches of ablating the projection offset $P_{o}$.
Learnable projection offset. In Setting II in Tab. 3, where the entire projection offset $P_{o}$ is eliminated, a significant drop in performance is observed. In this case, the model cannot successfully handle the occlusion effects as presented in Fig. 12 for both reconstruction and editing. In Setting III in Tab. 3, the removal of view-dependence from the learnable projection offset $P_{o}$ leads to a minor decrease in terms of statistical reconstruction, as the model no longer considers viewing directions.
5 Discussion and Conclusion
This paper explores the task of efficient 3D scene editing, where we focus on editing efficiency and user interactivity. Specifically, we propose AGAP, an editing-friendly and efficient solution for neural 3D scene editing, by representing a 3D scene using a 2D canonical image equipped with a projection field, such that users can easily perform efficient 3D editing via processing the 2D image. The key challenge and contribution of our method lies in how to regularize the projection field to make the canonical image look natural.
Compared to existing baselines, which require a time-consuming optimization process for one editing style, our approach can perform on-par 3D editing with at least 20$\times$ faster speed per edit and be far easier to use in practice. It supports various types of editing, such as scene stylization, instance segmentation, and interactive drawing.
Appendix A Supplementary Video
To offer a more comprehensive demonstration of our visual results, we have included a supplementary video showcasing three editing cases (scene stylization, instance segmentation, and texture editing) on diverse 3D scenes. Please check “demo.mp4” for details.
Appendix B Failure Cases and Limitation
Recall that in this paper, we come up with an editing-friendly representation, AGAP, which permits explicit 3D editing with the help of a natural 2D canonical image. In this section, we present some failure cases and discuss the limitation.
Our method supports texture editing by directly painting onto the canonical image. However, such painting might be distorted when the novel viewpoint exhibits occlusions on the edited regions. As shown in Fig. S1, we can easily paint an “AGAP” logo onto the marble pedestal of the fern plant in the canonical image, allowing us to directly obtain the edited NVS from different novel viewpoints through neural rendering. However, the logo appears distorted in the regions that are occluded by the fern plant.
Our pipeline includes a projection offset $P_{o}$ to handle moderate levels of occlusion, which implicitly projects and clusters the 3D points to nearby pixels. However, we acknowledge that our method has limitations in handling extensive occlusions in the 3D scenes. Suppose a 3D scene is extremely complex and contains numerously extensive occlusions. Projecting such a scene onto a 2D plane (like a UV map) is possible, but creating a 2D projection that naturally and fully displays the scene for easy interactivity is very challenging and nearly impossible. Hence, such cases are beyond the scope of our current study, and we mainly focus on 3D scenes with moderate levels of occlusion.
Appendix C Training Details
Our 3D editing pipeline involves a two-step process: (1) we first train a per-scene reconstruction model using the proposed AGAP representation, which includes an explicit density grid $\phi_{G}$, an explicit canonical image $\phi_{I}$, and an associated projection field $\phi_{P}$; (2) we can then perform explicit 2D edits on the canonical image $\phi_{I}$ for 3D scene editing, including scene stylization, instance segmentation, and texture editing. All experiments, including training on various scenes from different datasets, are conducted and tested on a single RTX A6000 GPU, with specific hyperparameter details outlined in Tab. S1.
Table S1: Hyperparameters for training various scenes in different datasets.
@@ Appendix C Training Details
  $\lambda_{uv}$ $\lambda_{tv}$
Forward-facing [45, 18] (768, *) $10^{-5}$ $10^{-5}$
Object-centric [46, 41, 24] (768, *) / (*, 768) $10^{-1}$ $10^{-4}$
Panorama [65, 17, 18, 3] (768, 1536) $10^{-1}$ $10^{-4}$
@@
$\phi_{G}^{*},\phi_{I}^{*},\phi_{P}^{*}=\operatorname*{arg\,min}_{\phi_{G},\phi_{I},\phi_{P}}{\mathcal{L}_{color}+\mathcal{L}_{uv}+\mathcal{L}_{tv}},$ (S1)
where the second and the third terms are controlled by their corresponding weight factors $\lambda_{uv}$ and $\lambda_{tv}$, respectively. To be specific, the weight factor $\lambda_{uv}$ is set as $10^{-5}$ for forward-facing scene and larger value of $10^{-1}$ or $10^{-2}$ for panorama and inward-facing 360 scenes; the weight factor $\lambda_{tv}$ is set as $10^{-4}$ for panorama scene and $10^{-5}$ for other scenes. Note that for panorama and inward-facing 360 data, the total variation term is disabled after $20000$ steps to learn depths in detail.
@@
At specific scaling-up milestone steps, we increase the $\phi_{G}$ voxel count by a factor of 2 and the $\phi_{I}$ pixel count by a factor of 4. For the forward-facing and object-centric data scenes, the voxel grid $\phi_{G}$ is scaled up at $\{2000,4000,6000,8000\}$ training steps and the canonical image $\phi_{I}$ is scaled up at $\{8000,16000\}$ training steps. For the panorama data types, the voxel grid $\phi_{G}$ is scaled up at $\{2000,4000,6000,8000,10000,12000,14000,16000\}$ training steps, and the canonical image $\phi_{I}$ is scaled up at $\{4000,8000,12000,16000\}$ training steps.
Size of voxel grid $\phi_{G}$ and canonical image $\phi_{I}$. After the progressive scaling up, The final resolution of the voxel grid $\phi_{G}$ is set as $384\times 384\times 256$ for forward-facing scenes and $320\times 320\times 320$ for other scenes.
For the NDC canonical camera of forward-facing scenes, we set the height $H_{I}$ of the learnable explicit canonical image $\phi_{I}$ as $768$, and the canonical image width $W_{I}$ is adaptively calculated according to the width-height aspect ratio of the training images and the computed bounding box of the scene in NDC space. Denoting the bounding box in NDC space as $(x^{\prime}_{min},x^{\prime}_{max})$ in $x^{\prime}$ dimension, $(y^{\prime}_{min},y^{\prime}_{max})$ in $y^{\prime}$ dimension, and $(z^{\prime}_{min},z^{\prime}_{max})=(-1,1)$ in $z^{\prime}$ dimension and the aspect ratio as $r_{I}$, we can then calculate the canonical image width as:
$W_{I}=H_{I}\times r_{I}\times\frac{x^{\prime}_{max}-x^{\prime}_{min}}{y^{\prime}_{max}-y^{\prime}_{min}}.$ (S2)
For the canonical camera of panorama scenes, the canonical image height is set to be $768$ and the width $W_{I}$ is set to be $2\times 768=1536$ according to the definition of Equirectangular projection.
For the canonical camera of object-centric scenes, the canonical image width and height are adaptive, where the canonical image width-height aspect ratio $r_{I}=\tfrac{W_{I}}{H_{I}}$ is calculated according to the uv range:
$r_{I}=\frac{u_{max}-u_{min}}{v_{max}-v_{min}},$ (S3)
where $u\in[-\pi,\pi]$ and $v\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]$. The shorter dimension, whether width or height, is set to $768$.